I need help with: Two (2) physics lab Questions on Ohms Law: Ohmic and Nonohmic Circuit Elements And three (3) lab questions on The ratio of e/ m for electrons.
Answer #1

1)    Here, I = 0.25 A

=> R' = 0.001 * 50 = 0.05 ohm

   => percent error = (0.001 * 50)/(0.2 * 200) * 100%

                           = 0.125 %

2)   percent error = (0.01 * 1)/(0.1 * 100) * 100%

                           = 0.1 %

1)   evB = mv2/r

   eV = 1/2mv2

Here, eliminating v2

=> e/m = 2V/(r2B2)

2)    speed of electrons = sqrt(2 * 1.6 * 10-19 * 60/9.11 * 10-31)

                                       = 4.59 * 106 m/sec

3)    Here,   B = (uo/2) * I/r

    =>   B = 2 * 3.14 * 10-7 * 2.5/0.01

              =   1.57 * 10-4 T

=> % error =    0.5 %